3. Consider the market for Sharpie markers (Q is a pack of 3 markers.) Suppose that...

  • - The demand curve for Sharpie is QD = 20 – P

  • - The supply curve for Sharpie is QS = 4P – 5

  • - The government imposes a tax of $1 per pack of Sharpies because of the foul smell

    (4 points each)

a. Before the tax is implemented, what is the equilibrium price paid by buyers, price received by sellers, and the number of Sharpie packs sold? P* = Pb = Ps = $5.00/pack; Q*= 15 packs

REV 2

  1. When the tax is implemented, what is the price paid by buyers, price received by sellers, and the number of Sharpie packs sold? Ps = 4.8, Pb = 5.8, Qtax = 14.2 packs

  2. What is the revenue generated from this tax? What is the deadweight loss (DWL) associated with the tax?

    Revenue=$14.2; DWL = $0.4

  3. What is the incidence of the tax to buyers? Use the price elasticity of demand and price elasticity of supply at the equilibrium point and the formula T = εS to find your answer (and show your work). Ed = -1*5/15 = -0.33,

b εS +|εD |

Es =4*5/15= 1.33; Tb = 0.8 (which matches finding from (b) that 80% of the $1 tax burden falls on buyers, who now pay $5.80 instead of $5)
Answer #1

Op = 20-P Qs = 4P-5 Before tax Op = Os 20-P = 4P-5 20+ 5 = 4P+P 25 = 58 &= 20-5 . pt = = 5 P6 = Ps 15 = After tax t = 1 Q = sPb = = Pstt 4.8 +1 5.8 = Quantity after tax Q> 4(4.82 - 5 = 9.2- 5 = 14.2 SS PO 5.8 -- p² 5 K - Ps 4.8 F - + - - DD 14.2 15 2Q = 20-P Ep = emba = LD - IED! ķ a= UP-5 Qs = do p = 4 ( 3 / 3 = 4 Buyers burden Es Es+ ED 413 w/vi/wht x 0.8 8001 -sellers burden = ED 0 Est ED a še WWW = 2 * 3 = 0.20 = 200).